# Think Smart, Act SMARL! Analyzing Probabilistic Logic Driven Safety in Multi-Agent Reinforcement Learning

### v1

@@ A.2. Centipede Game
Figure 9. Centipede game for 2 players with $p_{0}=2$ and growth rule $p_{t+1}\leftarrow p_{t}+2$. Red nodes represent player 1’s decision node and blue ones are player 2’s. The leaves represent the utilities received by each player. The first move is at the root of the tree.
A.2.1. Environment description
Prakash (2021) provides a good starting point with respect to implementation. However, in order to make the structure of the game homogeneous with the other environments experimented with [^1] This is not strictly required but makes the implementation of the Python environment, the interaction with agents, and the shield more straightforward., a simple modification was made that turns the centipede game into a repeated simultaneous game with changing utilities each round.
@@ A.2.1. Environment description
Figure 10. Repeated simultaneous game form of Centipede shown in Figure 9.
A.2.2. Experimental conditions
@@
A.3. 2-Player Extended Public Goods Game
@@
Figure 11. Expected payoff matrix for a 2-player EPGG with general endowments $c_{1}$ and $c_{2}$ (left) and the payoff when $c_{1}=c_{2}=2$ (right). For both, $\mu=\mathbb{E}_{f\sim\mathcal{N}(\mu,\sigma)}[f]$.
@@ A.3.1. Experimental conditions
A.4. Markov Stag-Hunt
Peysakhovich and Lerer (2018) describe a few grid environments that have the property that the optimal policies for agents reflect the structure of the Stag-Hunt game – principally, there are two strategies that the agents can converge to, either take on more risk by cooperating but attain higher rewards, or take on lower risk but get guaranteed small rewards. An optimal strategy may even be some mix of the two, depending on the details of the environment in question.
One of these environments, which they term the Markov Stag-Hunt environment has been adapted here to show that adding local probabilistic logic shields to MARL agents can help them converge to a global behavior that corresponds to either of the underlying pure Nash equilibria of choice – either the cooperative or the non-cooperative one. It should be stated that the original code for Markov Stag-Hunt from Peysakhovich and Lerer (2018) is not publicly available, and thus the environment created here is a custom approximation based on the sparse description in their paper. Thus, their results should not be directly compared ours.
A.4.1. Environment description
The adapted Markov Stag-Hunt environment used within this paper can be described in general as such: $n$ agents are placed in a square grid world of size $g\times g$. The agents have the following action set at each time step to move through the grid with respect to a global direction system: {UP, DOWN, LEFT, RIGHT, NONE}. There are $n_{stag}$ stags placed in the environment, as well as $n_{plant}$ plants, each occupying a single position each on the lattice. An exemplary visualisation of this environment can be seen in Figure 1(d). The agents can have either local or global observations, with or without a ‘self’ token, either as a 3D tensor or flattened one-hot encoded vector depending on the needs and configuration of the experiment.
@@ A.4.2. Experimental conditions
Appendix B Details on Shield Construction
B.1. Stag-Hunt
@@
For the mixed equilibria, we would need a more sophisticated shield, given the current limitations of the existing implementation of Yang et al. (2023). One way of doing this is to add soft constraints to the actions based on how far they differ from the mixed equilibrium. As a proxy for the policy of the agent, a set of the latest $h$ historical actions of the agents can be collected – this is termed a buffer. A mean policy $\hat{\pi}_{h}$ at the current time $t$ can then be derived from these actions, iterating over all $a^{\prime}\in\mathcal{A}$:
@@
For the experiments discussed below, the buffer length is set to $h=50$. Additionally, under the payoff matrix in Figure 8, the a priori mixed Nash equilibrium can be computed such that $\pi^{*}(\text{{Stag}}\mid s)=0.6$ and $\pi^{*}(\text{{Hare}}\mid s)=0.4$ for both agents. The evaluation safety (Figure 3(c)) is computed based on $\mathcal{T}_{pure}$ as it directly reflects the actions of the agents (with $\mathbf{P}_{\mathcal{T}_{pure}}(\text{{Stag}}\mid s)=1$ and $\mathbf{P}_{\mathcal{T}_{pure}}(\text{{Hare}}\mid s)=0$).
B.2. Centipede Game
@@
B.3. Extended Public Goods Game
@@
Let us instead create a shield that that uses a model $\hat{\mu}$ of the true $\mu$ estimated over training so far. Additionally, we can use the standard deviation of the multiplication factors so far, $\hat{\sigma}$ to model the uncertainty the agent has with respect to the current $f_{t}$ compared to $\hat{\mu}$. This can be done online without the need of a buffer using Welford’s method (Welford, 1962). We then compare $\hat{\mu}$ to the predetermined Nash outcomes and decide if it is high enough to warrant defecting – if $\hat{\mu}\geq 1$, cooperate, else defect. In Shield 5, this is set as the (Boolean) value of the fact sensor(mu_high) – its value is set to 1 if $\hat{\mu}\geq 1$ else 0. The agents will err on the side of a cooperative strategy when $21<\hat{\mu}<$ 2, where the rational strategy would be mixed.
@@
B.4. Markov Stag-Hunt
@@
Appendix C Single-Agent Experiment: CartSafe
@@
CartSafe is modification of OpenAI Gym’s classical control problem CartPole (Brockman et al., 2016) with simple constraints inspired by jemaw (2019). The problem has the following elements: the environment consists of a rod attached to a box or cart at one end and is allowed to freely swing around the attachment point. This setup exists in a bounded 2-dimensional real space with a set width and an active physics simulation. At each time step, the agent has two actions: accelerate the cart to the left or to the right. The agent observes the cart’s position, velocity, the angular velocity of the rod and the angle of the rod in radians. The goal is to actively balance the rod upright without leaving the bounds of the environment. The agent receives a reward at each time step based on the angle of the pole ($\theta_{pole}$) with respect to vertical: $r^{t}=1$ if $-15^{\circ}\leq\theta_{pole}\leq 15^{\circ}$ and $r^{t}=-1$ otherwise. The environment terminates when a maximum number of steps ($t_{max}$) are reached or the cart moves off screen. The additional constraints that are added to this setup are such that the cart must not enter the region bounded by the leftmost and rightmost quarters of the environment – thus the agent must learn to balance the rod near the center of the screen.
@@
C.1. Experimental conditions
The maximum length of each episode was set to $t_{max}=200$ with 500 training episodes. The agents are evaluated every 10 episodes. Each algorithm was re-run with 5 different seeds each and their aggregated results are discussed below. The PPO and DQN parameters were minimally hand-tuned and can be found in Table 2 in Appendix D. The PPO agents were shielded with safety penality coefficient $\alpha\in\{0.5,1.0,2.0\}$. The DQN-based algorithms were tested by varying the following aspects: on-policy vs. off-policy, $\epsilon$-greedy vs. softmax exploration (paired with greedy and softmax exploitation policies respectively) and whether they are shielded (SIQL with safety coefficient $\alpha\in\{1.0,5.0\}$) or not (IQL).
C.2. Shield construction
@@
Figure 12. Training, evaluation and safety results for CartSafe for PPO-based agents. The lines represent the mean and the shadow the standard deviation over 5 seeds. Results are smoothed using a rolling average with window size 50.
@@
C.3. Results: PLPG (Shielded PPO)
[…]
On-policy, $\epsilon$-greedy
Off-policy, softmax
On-policy, softmax
Figure 13. Training, evaluation and safety results for CartSafe for DQN-based agents with exploration policies $\epsilon$-greedy (evaluations executed with a greedy policy) and softmax (evaluations executed with a softmax policy) and on- and off-policy variations. The lines represent the mean and the shadow the standard deviation over 3 seeds.
C.4. Results: PLTD (Shielded DQN)
@@
Appendix D List of Hyperparameters
  • $\mathbf{\gamma}$: The discount factor.
  • lr/lr (actor)/lr (critic): Learning rate for Q network (DQN), actor/critic networks (PPO). All networks were optimised via Adam (Kingma and Ba, 2014), with all other hyperparameters left as in PyTorch’s default implementation (Paszke et al., 2019).
    $\epsilon\leftarrow\delta^{t}$ (9)
  • $\tau$: (DQN) Exploration factor for softmax exploration.
  • VF coef: (PPO) Value-function loss coefficient.
  • S coef: (PPO) Entropy loss coefficient.
  • $\epsilon$-clip: (PPO) Clipped policy update ratio range.
  • $\alpha$: Safety penalty coefficient (cf. Definition 1 and Yang et al. (2023)’s Definition 5.1).
@@ Appendix D List of Hyperparameters
Table 2. Hyperparameters for experiments.
Appendix E Full 2-Player EPGG Results
Figure 14 shows the results of agents playing the EPGG with multiplicative factor $f_{t}\sim\mathcal{N}(\mu,1)$ where $\mu\in\{0.5,1.5,2.5,5.0\}$. We see that in all cases, the IPPO agents improve their rewards over time, learning to play the instantaneous EPGG. The mean reward curves of the SIPPO agents either start out very close to the expected Nash strategy or quickly learn it.
We see that for $\mu=0.5$, the SIPPO agents show a much smaller variation in reward and behavior than the IPPO ones. The shielded agents have an immediate convergence towards the Nash strategy (non-cooperation) than the unshielded ones, although their rewards are also smaller. This is due to the fact that when $f_{t}\sim\mathcal{N}(0.5,1)$, there is a $\approx 30$% chance of the value being greater than 1, which, through cooperation, the agents can attain higher instantaneous rewards than if they play the expected strategy. However, since we wanted the agents to converge to the expected strategy rather than the instantaneous strategy, the fact that the agents obey the shield rather than the reward is a positive result. The low standard deviation for the SIPPO agents also imply that there is high certainty when they play Defect.
Likewise, when $\mu=5$, we see a similar trend – the SIPPO agents have a lower variance than the IPPO ones, but this time they converge quicker to the optimal strategy too, as they always cooperate. As $\mu$ is interpolated between 0.5 and 5, the optimal strategies flip from defect to cooperate. We see that for the shielded agents, the uncertainty whether to play this strategy increases around 1, where the agents play each action approximately equally, and decrease below and above 1, where it is more certain that a strategy is the expected Nash equilibrium over the episode. This results in a higher reward than the unshielded ones, with faster convergence. Except for $\mu=0.5$, the rewards for the SIPPO agents are either equal to or exceed the unshielded ones, and their cooperative or defective nature is much more predictable and faster to converge to.
$\mathbf{\mu=0.5}$
$\mathbf{\mu=1.0}$
$\mathbf{\mu=1.5}$
$\mathbf{\mu=2.5}$
$\mathbf{\mu=5.0}$
Figure 14. Training, evaluation and safety results for the 2-player Extended Public Goods Game for PLPG-based agents. The multiplicative factor $f_{t}$ is sampled from $\mathcal{N}(\mu,1)$, where $\mu\in\{0.5,1.0,1.5,2.5,5.0\}$ for different experiments. The lines represent the mean and the shadow the standard deviation over 5 seeds. Results are smoothed using a rolling average with window size 50.

### v2

@@ B.2 Centipede Game
Figure 8: Centipede game for 2 players with $p_{0}=2$ and growth rule $p_{t+1}\leftarrow p_{t}+2$. Red nodes represent player 1’s decision node and blue ones are player 2’s. The leaves represent the utilities received by each player. The first move is at the root of the tree.
B.2.1 Environment description
Prakash [25] provides a good starting point with respect to implementation. However, in order to make the structure of the game homogeneous with the other environments experimented with [^1] This is not strictly required but makes the implementation of the Python environment, the interaction with agents, and the shield more straightforward., a simple modification was made that turns the centipede game into a repeated simultaneous game with changing utilities each round.
@@ B.2.1 Environment description
Figure 9: Repeated simultaneous game form of Centipede shown in Figure 8.
B.2.2 Experimental conditions
@@
B.3 Two-Player Extended Public Goods Game
@@
Figure 10: Expected payoff matrix for a two-player EPGG with general endowments $c_{1}$ and $c_{2}$ (left) and the payoff when $c_{1}=c_{2}=2$ (right). For both, $\mu=\mathbb{E}_{f\sim\mathcal{N}(\mu,\sigma)}[f]$.
@@ B.3.1 Experimental conditions
B.4 Markov Stag-Hunt
Peysakhovich and Lerer [24] describe a few grid environments that have the property that the optimal policies for agents reflect the structure of the Stag-Hunt game – principally, there are two strategies that the agents can converge to, either take on more risk by cooperating but attain higher rewards, or take on lower risk but get guaranteed small rewards. An optimal strategy may even be some mix of the two, depending on the details of the environment in question.
One of these environments, which they term the Markov Stag-Hunt environment has been adapted here to show that adding local probabilistic logic shields to MARL agents can help them converge to a global behavior that corresponds to either of the underlying pure Nash equilibria of choice – either the cooperative or the non-cooperative one. It should be stated that the original code for Markov Stag-Hunt from [24] is not publicly available, and thus the environment created here is a custom approximation based on the sparse description in their paper. Thus, their results should not be directly compared ours.
B.4.1 Environment description
The adapted Markov Stag-Hunt environment used within this paper can be described in general as such: $n$ agents are placed in a square grid world of size $g\times g$. The agents have the following action set at each time step to move through the grid with respect to a global direction system: {UP, DOWN, LEFT, RIGHT, NONE}. There are $n_{stag}$ stags placed in the environment, as well as $n_{plant}$ plants, each occupying a single position each on the lattice. An exemplary visualisation of this environment can be seen in Figure 2. The agents can have either local or global observations, with or without a ‘self’ token, either as a 3D tensor or flattened one-hot encoded vector depending on the needs and configuration of the experiment.
@@ B.4.2 Experimental conditions
Appendix C Details on Shield Construction
C.1 Stag-Hunt
@@
For the mixed equilibria, we would need a more sophisticated shield, given the current limitations of the existing implementation of Yang et al. [39]. One way of doing this is to add soft constraints to the actions based on how far they differ from the mixed equilibrium. As a proxy for the policy of the agent, a set of the latest $h$ historical actions of the agents can be collected – this is termed a buffer. A mean policy $\hat{\pi}_{h}$ at the current time $t$ can then be derived from these actions, iterating over all $a^{\prime}\in\mathcal{A}$:
@@
For the experiments discussed below, the buffer length is set to $h=50$. Additionally, under the payoff matrix in Figure 7, the a priori mixed Nash equilibrium can be computed such that $\pi^{*}(\text{{Stag}}\mid s)=0.6$ and $\pi^{*}(\text{{Hare}}\mid s)=0.4$ for both agents. The evaluation safety is computed based on $\mathcal{T}_{pure}$ as it directly reflects the actions of the agents (with $\mathbf{P}_{\mathcal{T}_{pure}}(\text{{Stag}}\mid s)=1$ and $\mathbf{P}_{\mathcal{T}_{pure}}(\text{{Hare}}\mid s)=0$).
C.2 Centipede Game
@@
C.3 Extended Public Goods Game
@@
Let us instead create a shield that that uses a model $\hat{\mu}$ of the true $\mu$ estimated over training so far. Additionally, we can use the standard deviation of the multiplication factors so far, $\hat{\sigma}$ to model the uncertainty the agent has with respect to the current $f_{t}$ compared to $\hat{\mu}$. This can be done online without the need of a buffer using Welford’s method [37]. We then compare $\hat{\mu}$ to the predetermined Nash outcomes and decide if it is high enough to warrant defecting – if $\hat{\mu}\geq 1$, cooperate, else defect. In Shield 5, this is set as the (Boolean) value of the fact sensor(mu_high) – its value is set to 1 if $\hat{\mu}\geq 1$ else 0. The agents will err on the side of a cooperative strategy when $21<\hat{\mu}<$ 2, where the rational strategy would be mixed.
@@
C.4 Markov Stag-Hunt
@@
Appendix D Single-Agent Experiment: CartSafe
@@
CartSafe is modification of OpenAI Gym’s classical control problem CartPole [6] with simple constraints inspired by jemaw [18]. The problem has the following elements: the environment consists of a rod attached to a box or cart at one end and is allowed to freely swing around the attachment point. This setup exists in a bounded 2-dimensional real space with a set width and an active physics simulation. At each time step, the agent has two actions: accelerate the cart to the left or to the right. The agent observes the cart’s position, velocity, the angular velocity of the rod and the angle of the rod in radians. The goal is to actively balance the rod upright without leaving the bounds of the environment. The agent receives a reward at each time step based on the angle of the pole ($\theta_{pole}$) with respect to vertical: $r^{t}=1$ if $-15^{\circ}\leq\theta_{pole}\leq 15^{\circ}$ and $r^{t}=-1$ otherwise. The environment terminates when a maximum number of steps ($t_{max}$) are reached or the cart moves off screen. The additional constraints that are added to this setup are such that the cart must not enter the region bounded by the leftmost and rightmost quarters of the environment – thus the agent must learn to balance the rod near the center of the screen.
@@
D.1 Experimental conditions
The maximum length of each episode was set to $t_{max}=200$ with 500 training episodes. The agents are evaluated every 10 episodes. Each algorithm was re-run with 5 different seeds each and their aggregated results are discussed below. The PPO and DQN parameters were minimally hand-tuned and can be found in Table 4 in Appendix E. The PPO agents were shielded with safety penality coefficient $\alpha\in\{0.5,1.0,2.0\}$. The DQN-based algorithms were tested by varying the following aspects: on-policy vs. off-policy, $\epsilon$-greedy vs. softmax exploration (paired with greedy and softmax exploitation policies respectively) and whether they are shielded (SIQL with safety coefficient $\alpha\in\{1.0,5.0\}$) or not (IQL).
D.2 Shield construction
@@
Figure 11: Training, evaluation and safety results for CartSafe for on- and off-policy DQN-based agents with different exploration strategies: top row: off-policy, $\epsilon$-greedy; top middle row: on-policy, $\epsilon$-greedy; bottom middle row: off-policy, softmax; and bottom row: on-policy, softmax. The lines represent the mean and the shadow the standard deviation over 3 seeds.
@@
D.3 Results: PLTD (Shielded DQN)
@@
D.4 Results: PLPG (Shielded PPO)
Figure 12: Training, evaluation and safety results for CartSafe for PPO-based agents. The lines represent the mean and the shadow the standard deviation over 5 seeds. Results are smoothed using a rolling average with window size 50.
Appendix E List of Hyperparameters
@@
Table 4: Hyperparameters for experiments.
  • $\mathbf{\gamma}$: The discount factor.
  • lr/lr (actor)/lr (critic): Learning rate for Q network (DQN), actor/critic networks (PPO). All networks were optimised via Adam [19], with all other hyperparameters left as in PyTorch’s default implementation [23].
    $\epsilon\leftarrow\delta^{t}$ (18)
  • $\tau$: (DQN) Exploration factor for softmax exploration.
  • VF coef: (PPO) Value-function loss coefficient.
  • S coef: (PPO) Entropy loss coefficient.
  • $\epsilon$-clip: (PPO) Clipped policy update ratio range.
  • $\alpha$: Safety penalty coefficient (cf. Definition 3.3 and Yang et al. [39]’s Definition 5.1).
Appendix F Training Curves
The training curves for Stag-Hunt (Figure 13), 2-player Extended Public Goods Game (Figure 14) 2-player Markov Stag-Hunt (Figures 15 and 16) are given below. All remaining experiment training curves will be available via a Weights and Biases dashboard for convenience at the publication of this paper, or will be provided at the request of the reader.
Figure 13: Training, evaluation and safety results for Stag-Hunt for PLPG-based agents. Results are smoothed using a rolling average with window size 50. The lines represent the mean and the shadow the standard deviation over 5 seeds.
(a) Training ($\mu=0.5$)
(b) Evaluation ($\mu=0.5$)
(c) Safety ($\mu=0.5$)
(d) Training ($\mu=1.0$)
(e) Evaluation ($\mu=1.0$)
(f) Safety ($\mu=1.0$)
(g) Training ($\mu=1.5$)
(h) Evaluation ($\mu=1.5$)
(i) Safety ($\mu=1.5$)
(j) Training ($\mu=2.5$)
(k) Evaluation ($\mu=2.5$)
(l) Safety ($\mu=2.5$)
(m) Training ($\mu=5.0$)
(n) Evaluation ($\mu=5.0$)
(o) Safety ($\mu=5.0$)
Figure 14: Training, evaluation, and safety results for the two-player EPGG for PLPG-based agents. The multiplicative factor $f_{t}$ is sampled from $\mathcal{N}(\mu,1)$, where $\mu\in\{0.5,1.0,1.5,2.5,5.0\}$ for different experiments. The lines represent the mean and the shadow the standard deviation over 5 seeds. Results are smoothed using a rolling average with window size 50.
Figure 15: Behavioural and training results during training for Markov Stag-Hunt for PLPG-based agents with no shield or two different shields, $\mathcal{T}_{weak}$ or $\mathcal{T}_{strong}$. The lines represent the mean and the shadow the standard deviation over 3 seeds. Results are smoothed using a rolling average with window size 50.
Figure 16: Behavioural and training results for Markov Stag-Hunt for PLPG-based agents with or without parameter sharing during training, where one of the agents is shielded with $\mathcal{T}_{strong}$. The solid lines represent the mean and the shadow the standard deviation over 3 seeds. Results are smoothed using a rolling average with window size 50.